OpenStudy (anonymous):

Can someone help with this:

OpenStudy (anonymous):

wut do u need help on? post a pic or paste the question and maybe I can help :)

OpenStudy (anonymous):

It's a trig question: sin(4 * theta) = cos(2 * theta)

OpenStudy (anonymous):

ughhhh haven't learned that yet pal.. im sorry.. :(

OpenStudy (anonymous):

That's alright, thanks for trying!

OpenStudy (anonymous):

aww ok sorry again :)

OpenStudy (anonymous):

ima add u in case u need anything else

OpenStudy (anonymous):

there added u

OpenStudy (raden):

sin(4θ) = cos(2θ) 2sin(2θ)cos(2θ) = cos(2θ) 2sin(2θ)cos(2θ) - cos(2θ) = 0 factor out! cos(2θ) (2sin(2θ ) - 1) = 0 for the zeroes take each factor then setting be equal zero cos(2θ) = 0 or 2sin(2θ) - 1 = 0 solve for θ ...

OpenStudy (anonymous):

Thanks guys, really helped!

OpenStudy (raden):

you're welcome